how many pounds of coffee worth $8.00 should be added to 10 pounds of coffee worth $5.00 a pound to get a mixture worth $6.00 a pound
Found 2 solutions by stanbon, dabanfield:
Answer by stanbon(75887)

how many pounds of coffee worth $8.00 should be added to 10 pounds of coffee worth $5.00 a pound to get a mixture worth $6.00 a pound
-------
Equation:
value + value = value
8x + 5*10 = 6(x+10)
8x + 5*10 = 6x + 6*10
2x = 10
x = 5 lbs: (amt. of $8 coffee needed in the mixture)
==================
Cheers,
Stan H.

Answer by dabanfield(803)

how many pounds of coffee worth $8.00 should be added to 10 pounds of coffee worth $5.00 a pound to get a mixture worth $6.00 a pound
Let x be the amount of $8.00 coffee. Then when we are finished the total amount of $6.00 coffee we will have is 10 + x.
So:
10*5.00 + x*8.00 = (10+x)*6.00
50 + 8x = 60 + 6x
2x = 10
x = 5